# A Novel Technosignature Search in the Breakthrough Listen Green Bank Telescope Archive

### v1

@@ 3.4 Pipeline Overview
    Determines which blocks do not contain narrowband signals drifting at $<0.01$ Hz/s, and do not show signs of broadband or blips in the majority of their ON scans. These are deemed Level 2 blocks.
@@
    Inspect the remaining blocks visually to determine promising candidates.
The number of remaining candidates after each step can be found in Tables 5 and 6.
@@ 4 Results
We applied our detection algorithm to the bulk of the Breakthrough Listen Green Bank Telescope archive, consisting of $9,684$ cadences, each a collection of six 5-min observations. These observations spanned in frequency from L band to X band. Of this archive, 5,871 were ‘spliced’ files consisting of a observations across a range of nodes stitched together.
Our pipeline found a total of 23,688,878 Level 0 frequency blocks. Of these, 170,623 met our kurtosis criterion and were deemed Level 1 blocks. The distribution of the Level 0 blocks as a function of observing frequency can be seen in Figure 7. While some sections of the observing band have a higher number of frequency blocks than others, for the most part they are fairly evenly distributed across the frequency range. This is largely due to a pre-filtering of the especially dense RFI regions (primarily present in the L and S bands).
11,913 blocks survived filtering for broadband, blips, and zero-drift signals, and made it to Level 2. The distribution of the Level 2 blocks as a function of observing frequency can be seen in Figure 8. These 11,913 Level 2 blocks that passed all of our filters came from a total of 679 unique targets, and 106 of these targets contained more than 10 Level 2 blocks. Our pipeline picks up a wide variety of frequency blocks containing signals strong in the ON scans and weak in the OFF scans. Many of these signals do not display narrowband morphologies that traditional pipelines and machine learning approaches are designed to find. Figure 9 displays some examples of these kinds of signals. While not all of these signals are compelling candidates, they highlight the flexibility of our approach and the ability it has to pick up on a wide variety of signals. These might include signals with non-linear drift rates, signals that are not narrowband, or multiple signals in close proximity to each other.
We inspect each of our Level 2 blocks visually to determine which contain the most promising signals, which we label as our ‘candidates’. While this process could likely be automated to some degree in future work, as this is the first major application of this pipeline we do it manually to understand the kinds of signals that make it to Level 2.
@@
4.1 Most Promising Candidates
We find a total of 75 narrowband drifting signals that appear strongly in the ON scans of an observation and are not clearly present in the OFF scans. We highlight six of these candidates in Figure 10. These signals showed no sign of any signal in the OFF scans even after de-drifting them. These candidates serve as an example of what our follow up study entails for our candidates. Additional candidates can be found in Figure 16 and Figure 17 in Appendix A.
GJ 412B has a promising candidate at 1151.94082 MHz, and closer inspection reveals no sign of a signal in the OFF scans even after the observations were de-drifted and time-integrated. The closest other signal to this candidate is 38 MHz away, and is present in all scans. We deem this candidate signal likely to be RFI as a similar signal is found 12 kHz away from 1151.94082 MHz in observations of HIP 62207 taken a month earlier. This other signal drifts with exactly the same drift rate as our candidate, and is also significantly stronger in the ON scans than the OFF scans, although in this case it is weakly visible in the OFF scans. We do not expect that a technosignature would come from two different stars with exactly the same drift rate relative to Earth.
HIP 100064 and HIP 54677 also have promising signals, which did not appear at all in the OFF scans. These signals were also picked up by Price et al. (2020). They were deemed to be RFI due to the presence of similar signals with near identical drift rates at almost the same frequency in observations of other stars. We reach the same conclusion, and the candidate signals in both of these stars occur at exactly 1412.48393 MHz.
HIP 21556 has a promising signal at 1359.9441 MHz. The closest signal neighbor to this is 15 MHz away, and can be attributed to RFI as it appears strongly in the OFF scans. We deem this candidate as likely RFI due to a similar signal with near-identical drift rate is found 85 Hz away in observations of another target, HIP18280, which also weakens significantly in OFF scans.
HIP 99572 has promising signal at 1498.44515 MHz. This signal is in an RFI sparse environment, with no other signals present within 100 kHz. We deem it attributable to RFI due to similar signals appearing less than 10 kHz away in multiple other stars, including HIP 84012, HIP 54426, HIP 66207, and HIP 54677. All of these signals also are strong in the ON scans and weak in the OFF scans of their respective observations, and drift at similar rates.
@@ 4.1 Most Promising Candidates
We carry out a similar vetting process on all of our candidates, including those shown in Figures 16 and 17. Overall, we find that all of our most-promising candidates can likely be attributed to RFI.
@@ 4.2 Common False Positives
While we attribute all of the signals we pick up on to RFI, some are more clearly RFI than others. We highlight some of these common false positives in Figure LABEL:fig:false_pos.
Figures a) and b) display two signal morphologies that our pipeline struggles to vet out. The signal displayed in a) spans multiple frequency channels and is part of a larger structure. The signal in b) is a narrowband signal, with a non-linear drift. Both of these are signals that manifest in both ON and OFF scans, but weaken significantly in the OFF scans or strengthen significantly in the ON scans. As a result, two of the ON scans have very high kurtosis, while all of the OFF scans have low kurtosis. Neither of these signal get excised by our drift, broadband, or blip filters.
@@
Figures d) and e) show signals whose drift rates cause them to appear only in the ON scans of a given frequency block, while the signals themselves may appear strongly in the OFF scans outside of this frequency block. In figure d) this manifests as a high drift rate and non time-constant signal that can be seen extending into the OFF scans outside of the white dashed line which represent the frequency block our pipeline analyzes. In figure e) the signal’s drift rate is extremely variable, and it drifts in and out of the block’s frame.
@@ 5.1 Comparisons Across bands
L 455 4,433,776 60,995 4,588
S 710 3,834,698 91,903 2,984
C 3280 7,526,464 19,493 1,314
X 2600 8,754,368 17,801 2,347
Total 1.10–11.20 23,688,878 365,192 11,233
The highest block densities occur at L-band. This is primarily attributed to the higher number of satellites transmitting at 1.1 – 1.9 GHz. The density of Level 0 blocks is $\sim 3$ times as high as all other bands. The density of Level 2 blocks is more than 3 times as high as in S band, and $\sim 16$ times as high as the density of Level 2 blocks at C and X band. The fraction of Level 0 blocks that make it to Level 2 is $10.4\times 10^{-4}$ for L band, $7.7\times 10^{-4}$ for S band, $1.8\times 10^{-4}$ for C band, and $2.7\times 10^{-4}$ for X band. A much higher percentage of the RFI signals at L band weaken significantly in their respective OFF scans than in other bands. This may be due to the areal density and motion of satellites at L band, particularly GPS, having a higher tendency to cause RFI with a cadence around 10 minutes, and thus potentially appearing in successive ON scans.
@@
Level 0 Density 7.57 3.82 2.33 2.92
[[[[Blocks/Hz/Hr]
Level 1 Density 104.08 91.48 6.03 5.93
[[[[$10^{-3}$ Blocks/Hz/Hr]]]]
Level 2 Density 7.83 2.97 0.41 0.78
[[[[$10^{-3}$ Blocks/Hz/Hr]]]]
We also examine the spread in signal strength among the Level 0 and Level 2 blocks in each band. In L, S, and X band, the spread in signal-to-noise for signals in Level 0 blocks ranges from 5 up to $\sim 10^{3}$, whereas the signal-to-noise of signals only goes up to $\sim 3\times 10^{2}$ in C band. The spread of signal-to-noise for signals in Level 2 blocks follows the spread of Level 0 blocks closely in L and S band. In C and X band, the majority of Level 2 blocks have signals with SNR $<20$. We note that while we assume an SNR threshold of 10 when calculating the sensitivity of our search in Section 5.3, in practice our pipeline appears sensitive to signals with signal-to-noise ratios as low as $\sim 6$, as shown by Figure 13. However, we do not pick up on signals with SNR $<10$ with high completeness. We calculate the SNR of a signal by taking its mean over all time bins. While these signals need SNR $\geq 10$ in their eighth time bin in order to be detected by our pipeline, many of these signals are weaker in the other time bins of an ON scan, leading to an overall lower SNR for these signals.
@@ 5.3 Figures of Merit
It is difficult to quantify the thoroughness of a SETI search, due to the large parameter space covered and both known and unknown biases in target sampling. Various figures of merit (FoM) have been developed to assess the rigor of searches and allow for comparison between them. For calculations of all these FoMs in this section, we count only individual stars as targets, and do not include observations of many-star systems, such as globular clusters or galaxies. We also do not include the background fields encompassed by the GBT beam. As a result, the total stellar number is likely underestimated, although still a good representation of the number of targets we observe at the set distances we choose in Table 5.3.
@@
Figure 15: A comparison of transmitter rate vs. EIRPmin for this work with previous searches, where blue, red, green, and yellow stars correspond to searches of the L, S, C, and X bands respectively. The three vertical lines mark the EIRP of the Arecibo S-band radar, the power level of a Kardashev Type I civilization, and the total power budget of a Kardashev Type II civilization. The results of Enriquez et al. (2017), Price et al. (2020), Garrett & Siemion (2023), and Uno et al. (2023) are corrected for findings by Choza et al. (2024) that a turboseti search run with a SNR threshold of 10 in fact is only sensitive to signals with a SNR $\gtrsim 33$.
@@
where $k_{\rm B}$ is the Boltzmann constant, $T_{\rm sys}$ is the system temperature, and $A_{\rm eff}$ is the effective collecting area. The effective collecting area is a scaling of the physical collecting area of the telescope by some efficiency factor $\eta$ between 0 and 1. The SEFDs of GBT’s L, S, C, and X bands are reported in Table 5.3.
@@
Since our search used four different GBT receivers with varying fields of view, system temperatures, and bandwidths, we compute a combined $\rm{DFM}_{\rm tot}=\sum_{i}^{4}\rm{DFM}_{\rm i}$ which is simply the sum of each band’s DFM. A larger DFM in theory represents a more sensitive search. The $\rm{DFM}_{\rm tot}$ for this study is $8.39\times 10^{8}$ Hz Jy-3/2, $\times5$ as large as that of Price et al. 2020.
@@
Lower TFM scores represent more sensitive and more complete surveys. Our TFM values for each band are listed in Table 5.3. The total TFM score of this four-band survey is 6.7×6.7$\times$6.7 × smaller than the two-band search undertaken by Price et al. (2020).
@@
L band S band C band X band
TELESCOPE PARAMETERS
Antenna Diameter ( $D$ ) 100
Beam Width $(\theta)$ 8 4.5 1.7 1
Aperture Efficiency $(\eta)$ 0.72 0.68 0.65 0.63
@@
SEARCH PARAMETERS
Number of stars $<50$ pc 1134 1211 962 1123
Number of stars $<1000$ pc 2001 2124 1562 1826
Stellar Spectral Types All All All All
SNR Threshold 10 10 10 10
Spectral Resolution $(\delta\nu)[\mathrm{Hz}]$ 2.8 2.8 2.8 2.8
Band Frequency [GHz] $1.1-1.9$ $1.8-2.8$ $4.0-7.8$ $7.8-11.2$
Coverage [MHz] [of band] $455$ (57%) 710 (71%) 3280 (86%) 2600 (76%)
FIGURES OF MERIT
EIRPmin(50 pc) $2.05\times 10^{12}$ $2.32\times 10^{12}$ $2.76\times 10^{12}$ $3.07\times 10^{12}$
EIRPmin(1000 pc) $8.2\times 10^{14}$ $9.3\times 10^{14}$ $1.1\times 10^{15}$ $1.2\times 10^{15}$
CWTFM (50 pc) 0.43 0.32 0.26 0.28
DFM $3.05\times 10^{8}$ $2.49\times 10^{8}$ $1.85\times 10^{8}$ $1.0\times 10^{8}$
@@ 5.4 Transmitter Limit
We also calculate upper limits for transmitter strengths corresponding to a distance of 1000 pc. We use the EIRPmin values reported in Table 5.3. We find upper limits of $0.15\%$ for L band, $0.14\%$ for S band, $0.19\%$ for C band, and $0.16\%$ for X band.
@@ 5.5 Comparison to Previous Work
Multiple previous studies have analyzed GBT data as part of SETI surveys. The largest star-sample survey as of this work is Price et al. (2020), which analyzed 1,327 nearby stars across GBT’s L and S band, and the 10-cm receiver on the 64-m CSIRO Parkes telescope. Not all of the observations in Price et al. (2020)’s sample are also in ours. Specifically only 302 of their L band observations make it into our sample. However, all of the signals they highlight as promising come from stars whose observations we also analyze. In total, they showcase 8 of their best candidates, coming from 7 different stars. Of these 7 stars, our pipeline also identifies all of these sources as containing Level 2 blocks.
However, we are not able to pick up on all of their specific top ranked signals. Our pipeline only classifies 3 of their 8 signals as Level 2. For two of these candidates, this is because they share their frequency block with contaminating RFI (HIP100064 at 1595.87757 MHz and HIP1444 at 1449.98475 MHz). This increases the kurtosis of the OFF observations, and can also cause the frequency block to be thrown out if the RFI is not drifting. Another two signals (in HIP91699 at 2200.05036 MHz and HIP22845 at 2200.040568 MHz) occur at frequencies we deem RFI dense and do not analyze. Another signal (HIP109176 at 2414.52441 MHz) exhibits broadband behavior, which our algorithm filters out. While we do not pick up on all of the signals highlighted by Price et al. (2020), our pipeline does capture several dozen Level 2 blocks which contain equally if not more promising candidate signals. In fact, all of the signals shown in Figures 16 and 17 come from observations in Price et al. (2020)’s survey. This further highlights the need for novel search methods in picking up these promising candidate signals.
Ma et al. (2023) carried out a SETI search on 820 stars observed with the GBT L-band. They used a $\beta$-Convolutional Variational Autoencoder as opposed to traditional pipelines such as turboseti. They trained their autoencoder on synthetic signals generated with setigen. They identified eight promising candidate signals from five different targets. All five of these targets are picked up on by our pipeline as containing Level 2 blocks in their observations.
@@
While we do not recover $100\%$ of the signals highlighted by previous searches, we find many signals of comparable quality from the same sources. In addition, we believe that our pipeline would be successful at finding many of these signals with a slightly modified set of tuning parameters. Due to the size of our search and the quantity of data we analyzed, we imposed strict restrictions on the strength of the signal, the drift rates we searched over, and the the frequency ranges we chose to cut out. Our current false positive rates are $<1\%$, and weaker signals could be picked up on at the expense of a slightly higher false positive rate if the $k_{\rm ON,med}$ cutoff is lowered. Additionally, if a smaller number of observations is being analyzed, a smaller fraction of the frequency coverage could be blocked out. The primary reason we remove it is to speed up computation time for the large number of files we process. We find that we successfully recover the signals found by Price et al. (2020) and Ma et al. (2023) in regions we deem RFI dense if we do not cut these regions out preliminarily. However, we also note that a large fraction of these signals that occur in RFI dense regions deserve to be cut. For example, two of Price et al. (2020)’s candidates occur at 2200 MHz, a common frequency for satellite communications. Additionally, our pipeline uncovers additional high-quality candidate signals in the dataset used by Price et al. (2020) that they did not find using turboSETI.

### v2

@@ 3.4 Pipeline Overview
    Determines which blocks do not contain narrowband signals drifting at <0.01<0.01 Hz/s, and do not show signs of broadband or blips in the majority of their ON scans. These are deemed Level 2 blocks.
@@
    Inspect the remaining blocks visually to determine final candidates for further analysis.
The number of remaining frequency blocks after each step can be found in Tables 5 and 6.
@@ 4 Results
We applied our detection algorithm to the bulk of the Breakthrough Listen Green Bank Telescope archive, consisting of $6,630$ cadences, each a collection of six 5-min observations. These observations spanned in frequency from L band to X band. Of this archive, 2,138 were ‘spliced’ files consisting of a observations across a range of nodes stitched together.
Our pipeline found a total of $24,368,715$ Level 0 frequency blocks. Of these, $175,160$ met our kurtosis criterion and were deemed Level 1 blocks. The distribution of the Level 0 blocks as a function of observing frequency can be seen in Figure 7. While some sections of the observing band have a higher number of frequency blocks than others, for the most part they are fairly evenly distributed across the frequency range. This is largely due to a pre-filtering of the especially dense RFI regions (primarily present in the L and S bands).
$14,168$ blocks survived filtering for broadband, blips, and zero-drift signals, and made it to Level 2. The distribution of the Level 2 blocks as a function of observing frequency can be seen in Figure 8. These $14,168$ Level 2 blocks that passed all of our filters came from a total of 838 unique targets, and 126 of these targets contained more than 10 Level 2 blocks. Our pipeline picks up a wide variety of frequency blocks containing signals strong in the ON scans and weak in the OFF scans. Many of these signals do not display narrowband morphologies that traditional pipelines and machine learning approaches are designed to find. Figure 9 displays some examples of these kinds of signals. While not all of these signals are compelling candidates, they highlight the flexibility of our approach and the ability it has to pick up on a wide variety of signals. These might include signals with non-linear drift rates, signals that are not narrowband, or multiple signals in close proximity to each other.
We inspect each of our Level 2 blocks visually to determine which contain signals worthy of further analysis, which we label as our ‘candidates’. While this process could likely be automated to some degree in future work, as this is the first major application of this pipeline we do it manually to understand the kinds of signals that make it to Level 2.
@@
4.1 Candidates Surviving Visual Inspection
We find a total of 75 narrowband drifting signals that appear strongly in the ON scans of an observation and are not clearly present in the OFF scans. We highlight six of these candidates in Figure 10. These signals showed no sign of any signal in the OFF scans even after de-drifting them. These candidates serve as an example of what our follow up study entails for our candidates. Additional candidates can be found in Figure 21 and Figure 22 in Appendix B.
GJ 412B has a candidate signal at 1151.94082 MHz, and closer inspection reveals no sign of a signal in the OFF scans even after the observations were de-drifted and time-integrated. The closest other signal to this candidate is 38 MHz away, and is present in all scans. We deem this candidate signal likely to be RFI as a similar signal is found 12 kHz away from 1151.94082 MHz in observations of HIP 62207 taken a month earlier. This other signal drifts with exactly the same drift rate as our candidate, and is also significantly stronger in the ON scans than the OFF scans, although in this case it is weakly visible in the OFF scans. We do not expect that a technosignature would come from two different stars with exactly the same drift rate relative to Earth.
HIP 100064 and HIP 54677 also have candidate signals, which did not appear at all in the OFF scans. These signals were also picked up by Price et al. (2020). They were deemed to be RFI due to the presence of similar signals with near identical drift rates at almost the same frequency in observations of other stars. We reach the same conclusion, and the candidate signals in both of these stars occur at exactly 1412.48393 MHz.
HIP 21556 has a candidate signal at 1359.9441 MHz. The closest signal neighbor to this is 15 MHz away, and can be attributed to RFI as it appears strongly in the OFF scans. We deem this candidate as likely RFI due to a similar signal with near-identical drift rate is found 85 Hz away in observations of another target, HIP18280, which also weakens significantly in OFF scans.
HIP 99572 has a candidate signal at 1498.44515 MHz. This signal is in an RFI sparse environment, with no other signals present within 100 kHz. We deem it attributable to RFI due to similar signals appearing less than 10 kHz away in multiple other stars, including HIP 84012, HIP 54426, HIP 66207, and HIP 54677. All of these signals also are strong in the ON scans and weak in the OFF scans of their respective observations, and drift at similar rates.
@@ 4.1 Candidates Surviving Visual Inspection
We carry out a similar vetting process on all of our candidates, including those shown in Figures 21 and 22. Overall, we find that all of our candidates can likely be attributed to RFI.
@@ 4.2 Common False Positives
While we attribute all of the signals we pick up on to RFI, some are more clearly RFI than others. We highlight some of these common false positives in Figure 11.
Figures a) and b) display two signal morphologies that our pipeline struggles to vet out. The signal displayed in a) spans multiple frequency channels and is part of a larger structure. The signal in b) is a narrowband signal, with a non-linear drift. Both of these are signals that manifest in both ON and OFF scans, but weaken significantly in the OFF scans or strengthen significantly in the ON scans. As a result, two of the ON scans have very high kurtosis, while all of the OFF scans have low kurtosis. Neither of these signals is excised by our drift, broadband, or blip filters.
@@
Figures d) and e) show signals whose drift rates cause them to appear only in the ON scans of a given frequency block, while the signals themselves may appear strongly in the OFF scans outside of this frequency block. In figure d) this manifests as a high drift rate and non time-constant signal that can be seen extending into the OFF scans outside of the white dashed line which represent the frequency block our pipeline analyzes. In figure e) the signal’s drift rate is extremely variable, and it drifts in and out of the block.
@@ 5.1 Comparisons Across bands
L 455 4,196,680 101,568 6,131
S 710 3,392,299 33,308 3,382
C 3280 8,423,757 21,751 2,838
X 2600 8,355,979 18,533 1,817
Total 1.10–11.20 24,368,715 175,160 14,168
The highest block densities occur at L-band. This is primarily attributed to the higher number of satellites transmitting at 1.1 – 1.9 GHz. The density of Level 1 blocks is more than 3 times as high as in S band, and $\sim 20$ times as high as the density of Level 2 blocks at C and X band. The fraction of Level 0 blocks that make it to Level 2 is $1.46\times 10^{-3}$ for L band, $1.0\times 10^{-3}$ for S band, $3.4\times 10^{-4}$ for C band, and $2.1\times 10^{-4}$ for X band. A much higher percentage of the RFI signals at L band weaken significantly in their respective OFF scans than in other bands. This may be due to the areal density and motion of satellites at L band, particularly GPS, having a higher tendency to cause RFI with a cadence around 10 minutes, and thus potentially appearing in successive ON scans.
@@
Level 0 Density 8.7 6.3 4.6 4.1
[[Blocks/Hz/Hr]
Level 1 Density 209.4 61.6 11.8 9.2
$10^{-3}$ Blocks/Hz/Hr]]
Level 2 Density 12.6 6.3 1.5 0.9
$10^{-3}$ Blocks/Hz/Hr]]
We also examine the spread in signal strength among the Level 0 and Level 2 blocks in each band. In L, S, and X band, the spread in signal-to-noise for signals in Level 0 blocks ranges from 5 up to $\sim 10^{3}$, whereas the signal-to-noise of signals only goes up to $\sim 3\times 10^{2}$ in C band. The spread of signal-to-noise for signals in Level 2 blocks follows the spread of Level 0 blocks closely in L and S band. In C and X band, the majority of Level 2 blocks have signals with SNR <20<20. We note that while we assume an SNR threshold of 10 when calculating the sensitivity of our search in Section 5.3, in practice our pipeline appears sensitive to signals with signal-to-noise ratios as low as $\sim 6$, as shown by Figure 13. However, we do not pick up on signals with SNR <10<10 with high completeness. We calculate the SNR of a signal by taking its mean over all time bins. While these signals need SNR $\geq 10$ in their eighth time bin in order to be detected by our pipeline, many of these signals are weaker in the other time bins of an ON scan, leading to an overall lower SNR for these signals.
@@ 5.2 Parameter Space Probed and Common Signal Morphologies
We also examine the distribution of the pointing directions of the observations containing our candidates and Level 2 blocks. We plot the altitude and azimuth of of these pointings in Appendix A, and find that many of our Level 2 blocks originate from L band observations of sources located at an azimuth of $210\arcdeg\pm 35\arcdeg$. A large fraction of our candidates also originate from this region. This could correspond to a large cluster of geosynchronous satellites as viewed from the GBT. We carry out a search to match our candidate signals to known satellites whose paths pass near to the observation pointing, but are unable to succesfully confirm that any of our candidate signals come from specific satellites.
@@ 5.3 Figures of Merit
It is difficult to quantify the thoroughness of a SETI search, due to the large parameter space covered and both known and unknown biases in target sampling. Various figures of merit (FoM) have been developed to assess the rigor of searches and allow for comparison between them. For calculations of all these FoMs in this section, we count only individual stars as targets, and do not include observations of many-star systems, such as globular clusters or galaxies. We also do not include the background fields encompassed by the GBT beam. As a result, the total stellar number is likely underestimated, although still a good representation of the number of targets we observe at the set distances we choose in Table 7.
@@
Figure 15: A comparison of transmitter rate vs. EIRPmin for this work with previous searches, where blue, red, green, and yellow stars correspond to searches of the L, S, C, and X bands respectively. The three vertical lines mark the EIRP of the Arecibo S-band radar, the power level of a Kardashev Type I civilization, and the total power budget of a Kardashev Type II civilization. The results of Enriquez et al. (2017), Price et al. (2020), Garrett & Siemion (2023), and Uno et al. (2023) are corrected for findings by Choza et al. (2024) that a turboseti search run with a SNR threshold of 10 in fact is only sensitive to signals with a SNR $\gtrsim 33$. Other results shown come from Tremblay & Tingay (2020), Harp et al. (2016), Siemion et al. (2013), Backus & Project Phoenix Team (2004), Horowitz & Sagan (1993).
@@
where $k_{\rm B}$ is the Boltzmann constant, $T_{\rm sys}$ is the system temperature, and $A_{\rm eff}$ is the effective collecting area. The effective collecting area is a scaling of the physical collecting area of the telescope by some efficiency factor $\eta$ between 0 and 1. The SEFDs of GBT’s L, S, C, and X bands are reported in Table 7.
@@
Since our search used four different GBT receivers with varying fields of view, system temperatures, and bandwidths, we compute a combined $\rm{DFM}_{\rm tot}=\sum_{i}^{4}\rm{DFM}_{\rm i}$ which is simply the sum of each band’s DFM. A larger DFM in theory represents a more sensitive search. The $\rm{DFM}_{\rm tot}$ for this study is $4.8\times 10^{8}$ Hz Jy-3/2, $3.1\times$ as large as that of Price et al. 2020.
@@
Lower TFM scores represent more sensitive and more complete surveys. Our TFM values for each band are listed in Table 7. The total TFM score of this four-band survey is $3.2\times$ smaller than the two-band search undertaken by Price et al. (2020).
@@ 5.4 Transmitter Limit
We also calculate upper limits for transmitter strengths corresponding to a distance of 1000 pc. We use the EIRPmin values reported in Table 7. We find upper limits of $0.20\%$ for L band, $0.29\%$ for S band, $0.37\%$ for C band, and $0.24\%$ for X band.
@@ 5.5 Comparison to Previous Work
Multiple previous studies have analyzed GBT data as part of SETI surveys. The largest star-sample survey as of this work is Price et al. (2020), which analyzed 1,327 nearby stars across GBT’s L and S band, and the 10-cm receiver on the 64-m CSIRO Parkes telescope. Not all of the observations in Price et al. (2020)’s sample are also in ours. Specifically only 981 of their L band observations make it into our sample. However, all of the signals they highlight as interesting come from stars whose observations we also analyze. In total, they showcase 8 of their best candidates, coming from 7 different stars. Of these 7 stars, our pipeline also identifies all of these sources as containing Level 2 blocks.
However, we are not able to pick up on all of their specific top ranked signals. Our pipeline only classifies 3 of their 8 signals as Level 2. For two of these candidates, this is because they share their frequency block with contaminating RFI (HIP100064 at 1595.87757 MHz and HIP1444 at 1449.98475 MHz). This increases the kurtosis of the OFF observations, and can also cause the frequency block to be thrown out if the RFI is not drifting. Another two signals (in HIP91699 at 2200.05036 MHz and HIP22845 at 2200.040568 MHz) occur at frequencies we deem RFI dense and do not analyze. Another signal (HIP109176 at 2414.52441 MHz) exhibits broadband behavior, which our algorithm filters out. While we do not pick up on all of the signals highlighted by Price et al. (2020), our pipeline does capture several dozen Level 2 blocks which contain equally if not more interesting candidate signals. In fact, all of the signals shown in Figures 21 and 22 come from observations in Price et al. (2020)’s survey. This further highlights the need for novel search methods in picking up strong candidate signals.
Ma et al. (2023) carried out a SETI search on 820 stars observed with the GBT L-band. They used a $\beta$-Convolutional Variational Autoencoder as opposed to traditional pipelines such as turboseti. They trained their autoencoder on synthetic signals generated with setigen. They identified eight candidate signals from five different targets. All five of these targets are picked up on by our pipeline as containing Level 2 blocks in their observations.
@@
While we do not recover $100\%$ of the signals highlighted by previous searches, we find many signals of comparable quality from the same sources. In addition, we believe that our pipeline would be successful at finding many of these signals with a slightly modified set of tuning parameters. Due to the size of our search and the quantity of data we analyzed, we imposed strict restrictions on the strength of the signal, the drift rates we searched over, and the the frequency ranges we chose to cut out. Our current false positive rates are $1\%$, and weaker signals could be picked up on at the expense of a slightly higher false positive rate if the $k_{\rm ON,med}$ cutoff is lowered. Additionally, if a smaller number of observations is being analyzed, a smaller fraction of the frequency coverage could be blocked out. The primary reason we remove it is to speed up computation time for the large number of files we process. We find that we successfully recover the signals found by Price et al. (2020) and Ma et al. (2023) in regions we deem RFI dense if we do not cut these regions out preliminarily. However, we also note that a large fraction of these signals that occur in RFI dense regions deserve to be cut. For example, two of Price et al. (2020)’s candidates occur at 2200 MHz, a common frequency for satellite communications. Additionally, our pipeline uncovers additional high-quality candidate signals in the dataset used by Price et al. (2020) that they did not find using turboSETI.